A tubular-bowl centrifuge is used to concentrate a suspension of genetically-engineered yeast containing a new recombinant protein. At a speed of 12000 rpm, the centrifuge treats broth of 3 L/min with satisfactory results. It is proposed to use the same centrifuge to separate cell debris from homogenate produced by mechanical disruption of the yeast. If the average size of the debris is one-third of the yeast and the viscosity of the homogenate is five times greater than the cell suspension, what flow rate can be handled if the centrifuge is operated at the same speed?


1
Expert's answer
2021-01-19T04:31:12-0500

find the foкmula:

"D=\frac{W}{V}"

D - the dilution factor

"D=\frac{1}{3}=0.333"

"0.333=\frac{W}{1}"

W=0.333 m3 / min